Circular No. 7569

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001G IN IC 677 A. Gal-Yam and O. Shemmer, Tel Aviv University, report: “A spec-

trum (range 400–780 nm), obtained with the Wise Observatory 1-m telescope (+ FOSC spectrograph) on Jan. 16 by J. Dann, reveals that SN 2001G (cf. IAUC 7560) is of type Ia, near maximum brightness. We estimate a redshift of 0.015 from the supernova’s broad features.”

SUPERNOVA 2001N IN NGC 3327 S. Jha, T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001N (cf. IAUC 7568), obtained by M. Calkins on Jan. 22.46 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova well before maximum light. Using the NASA/IPAC Extragalactic Database recession velocity of 6303 km/s for the host galaxy, the supernova expansion velocities are 13 300 and 20 600 km/s for Si II (rest 635.5 nm) and Ca II H and K, respectively. These high velocities imply that the supernova is ∼ 10 days before maximum light (Jha et al. 1999, Ap.J. Supp. 125, 73).

SUPERNOVA 2001I IN UGC 2836 A. Dimai, Cortina d’Ampezzo, Italy, reports that an image taken on

Jan. 15.75 UT with a 0.5-m telescope shows SN 2001I at mag ∼ 17.5 (independent discovery); an image taken on 2000 Dec. 22.8, however, showed nothing at the position of SN 2001I (cf. IAUC 7563) to limiting mag 18.3.

SUPERNOVA 2001K IN IC 677 G. M. Hurst, Basingstoke, England, reports the independent discovery

of SN 2001K (cf. IAUC 7565) by M. Armstrong, Rolvenden, Kent, who provides the following unfiltered CCD magnitude estimates (using USNO-A2.0 red magnitudes): 1999 Dec. 2 UT, [18.5; 2001 Jan. 16.203, 16.6; 18.029, 16.3. He also provides the following position end figures (average from measures of two CCD frames): 56^{s}.14, 05′′.6.

V445 PUPPIS Photometry by A. C. Gilmore (cf. IAUC 7559): Jan. 11.50 UT, V =

8.81, U–B = −0.43, B–V = +0.61, V –R = +0.41, V –I = +1.00; 18.45, 9.67, −0.37, +0.61, +0.45, +1.10; 19.45, 9.40, −0.41, +0.62, +0.43, +1.05.

2001 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green